OpenStudy (chosenmatt):

What is the approximate area of a circle with a radius of 13 cm?

OpenStudy (chosenmatt):

@iGreen

OpenStudy (chosenmatt):

@iGreen

OpenStudy (igreen):

\(A = \pi r^2\) \(A = (3.14)(13^2)\) Simplify

OpenStudy (chosenmatt):

40.8 530.7 784 2461.8

OpenStudy (chosenmatt):

^^answer choices @iGreen

OpenStudy (chosenmatt):

@iGreen @satellite73

OpenStudy (igreen):

Multiply 3.14 * 13 * 13

OpenStudy (chosenmatt):

530.66

OpenStudy (igreen):

Yes, that rounds to 530.7

OpenStudy (chosenmatt):

oh